Find the equation of the right circular cone whose vertex is (1,−1,2), the axis is

(x−1)/2 =( y+1)/1 = (z−2)/-2

and the semi-vertical angle is 45◦ .



1
Expert's answer
2021-06-07T18:43:38-0400

Let "P(x, y, z)" be any point on the cone. The direction ratio from the origin(axis) to point on the cone is "x-1, y+1, z-2" and the direction ratio of the axis is "2, 1, -2."

Thus, 


"\cos 45\degree=\dfrac{2(x-1)+1(y+1)-2(z-2)}{\sqrt{(x-1)^2+(y+1)^2+(z-2)^2}\sqrt{(2)^2+(1)^2+(-2)^2}}"


"\dfrac{1}{\sqrt{2}}=\dfrac{2x+y-2z+3}{3\sqrt{(x-1)^2+(y+1)^2+(z-2)^2}}"

On simplifying this we get


"9x^2-18x+9+9y^2+18y+9+9z^2-36z+36"

"=8xy-16xz+8x^2+24x+2y^2+12y-8yz"

"+8z^2-24z+18"

"x^2+7y^2+z^2-8xy+16xz+8yz-42x+6y-12z+36=0"